1. f the price of a book is reduced by Rs.5 a person can buy 5 more books for Rs.300. Find the original price of the book.

Grade:10

6 Answers

Devesh Lall
12 years ago
SOLUTION FOR QUESTION NO.1
Let the original price of the book be x
Since if the price is reduced by Rs.5, then let the new price be Rs.(x-5)
If 5 more books are to be bought then req. price will be =Rs5(x-5)
Required price given=Rs.300
Now,
A/Q:-
5(x-5) = 300
x = Rs.65
PHANI TEJA
12 years ago
let the original price be x
reduced price=x-5
cost of 5 books after reducing the price=5(x-5)
therefore,
5(x-5)=300
x-5=60
x=65
Kishor Mohite
12 years ago
let initially no. of books person can buy = n
price of each book = x
person has 300 Rs hence,
nx=300  ...(1)
if price of each book is reduced by Rs. 5 he can buy 5 more books in 300 Rs.
(n+5)(x-5)=300
nx-5n+5x-25=300  ...(2)
from eq. 1 nx=300 and n=300/x
so eq. 2 reduces to
x2-5x-300=0
so x=20 or -15
hence initial price of each book = 20 Rs.

Khizar Iqbal
7 years ago
f the price of a book is reduced by Rs.5 a person can buy 5 more books for Rs.300. Find the original price of the book.
The person has purchased 5 more books ie the has purchased Total 6 books
therefore
in simple 300/6 = 50 Rs. Earlier before reducing 5 Rs the price of the each book was 55 Rs.
Anuj gupta
6 years ago
Xy=300Y=300/x(X-5)(y+5)=300Xy+5x-5y -25=300300+5x-1500/x-25=3005x2-25x-1500=0X2-5x-300=0X2+15x-20x-300=0X(x+15)-20(x+15)(X-20)(x+15)X=20,-15X=20
Deeva
6 years ago
300/x–5 – 300/x = 5300x–300x+1500/xx–5x= 5300/xx–5x=0Eq=xx–5x–300Solving by qua formala√d=35X=20 and X=–15–15 is rejectedSo X=20rs